### v1

@@ Abstract
There is growing interest in automating surgical tasks using robotic systems, such as endoscopy for treating gastrointestinal (GI) cancer. However, previous studies have primarily focused on detecting and analyzing objects or robots, with limited attention to ensuring safety, which is critical for clinical applications, where accidents can be caused by unsafe robot motions. In this study, we propose a new control framework that can formally ensure the safety of automating certain processes involved in endoscopic submucosal dissection (ESD), a representative endoscopic surgical method for the treatment of early GI cancer, by using an endoscopic robot. The proposed framework utilizes Control Barrier Functions (CBFs) to accurately identify the boundaries of individual tumors, even in close proximity within the GI tract, ensuring precise treatment and removal while preserving the surrounding normal tissue. Additionally, by adopting a model-free control scheme, safety assurance is made possible even in endoscopic robotic systems where dynamic modeling is challenging. We demonstrate the proposed framework in cases where the tumors to be removed are close to each other, showing that the safety constraints are enforced. We show that the model-free CBF-based controlled robot eliminates one tumor completely without damaging it, while not invading another nearby tumor.
@@ I Introduction
Virtual fixture is a method that defines virtual boundaries or constraints to guide the movements of a robot along a specific trajectory or restrict its operation within a designated space [30]. Especially in the field of surgical robotics, it has been widely applied to improve safety by preventing the robot from approaching sensitive organs or lesions [31]. By using both guidance virtual fixtures and forbidden region virtual fixtures together, the surgical robot can track an accurate trajectory while avoiding areas that should not be accessed. However, the mathematical expressions that define virtual fixture constraints typically exhibit discontinuities. This is an issue, because such discontinuous constraints can cause damage to the lesion tissue due to unintended movements, significantly impacting the quality and reliability of robotic surgery. Furthermore, virtual fixture-based approaches lack formal safety guarantees, which might lead to serious safety violations.
@@
Based on Molnar’s work [40], we propose a new framework that can formally ensure the safety of certain processes involved in endoscopic submucosal dissection (ESD), a representative endoscopic surgical method for the treatment of early GI cancer [1, 2, 42], by using an endoscopic robot (Fig. 1). The purpose of ESD is not only to completely remove the tumor from the GI tract but also to ensure that the entire tumor is resected as a single piece, allowing for accurate pathological examination. In summary, as shown in Fig. 2, the framework proposed in this study enables the automated motion of the robot to clearly distinguish the boundaries of each tumor, even when one or more tumors are located close to each other within the GI tract. This allows for the individual removal of each tumor while preserving the normal mucosal areas between tumors.
The main contribution of this study can be summarized as follows: This paper proposes an integrated framework that includes a model-free safety-critical control scheme to ensure safety in ESD robotic automation. The framework plans safe motions for tumor removal without affecting surrounding tumors, defining safety boundaries for each tumor and enforcing them as constraints in a model-free CBF-based safe controller to ensure safety.
@@ II-C Input-to-State Stability
where the bounded disturbance, $\bm{d}\in\mathbb{R}^{m}$, deteriorates the performance of CLF in terms of stability. To resolve the impact of disturbances, the concept of Input-to-State exponential Stability (ISS) [43] can be introduced in (3) as follows:
@@
where $\iota$ is an extended class $\mathcal{K}^{e}_{\infty}$ function; thus, we can achieve a robust stable controller satisfying (7). It is worth noting that the concept allows the controller to converge to a neighborhood of the origin, is case of a nonzero disturbance.
@@ II-E Model-Free Robotic Safety-Critical Control
There exists the exponentially stable velocity controller, $\bm{u}=k_{q}(\bm{q},\dot{\bm{q}})$ where $k_{q}:\mathrm{Q}\times\mathbb{R}^{n}\rightarrow\mathcal{U}$, satisfying $||\dot{\bm{e}}(t)||\leq M||\dot{\bm{e}}_{0}||\mathrm{e}^{-\lambda t}$ for the positive scalar, $M,\lambda$, where $\dot{\bm{e}}=\dot{\bm{q}}-\dot{\bm{q}}_{d}$ is the tracking error between current velocity and desired one. In other words, a continuously differentiable Lyapunov function, $V(\bm{q},\dot{\bm{q}})$ exists such that:
$k_{1}||\dot{\bm{e}}||\leq V(\bm{q},\dot{\bm{e}})\leq k_{2}||\dot{\bm{e}}||,$ (12)
@@
and a stable velocity tracking controller satisfying (13). If the velocity controller tracks a safe velocity, $\dot{\bm{q}}_{s}$, fast enough, and if $\lambda>\alpha$ such that the initial conditions, $(\bm{q}_{0},\dot{\bm{e}}_{0})\in\mathcal{S}_{V}$ can be defined as:
@@
Figure 2: Illustration of the proposed framework for safe robotic ESD automation. Once the tumor is detected by the robot, the safe set boundary is defined to ensure safety. Then, a series of marking points are automatically generated around the tumor. Finally, cutting is performed along the trajectory defined by these marking points using a CBF-based safety-critical velocity controller.
III Application to Safety-Critical Robotic ESD Automation
Fig. 2 illustrates the framework we propose for ensuring the formal safety in ESD automation. ESD can be subdivided into four main processes: injection, marking, cutting, and dissection [46]. Considering the difficulty of each process involved in ESD and the potential clinical safety issues, this study focuses on developing a method to ensure the safety of marking and cutting automation.
@@ III-A Automatic Definition of Marking Points
The marking process is to mark specific points around each tumor. Since the boundaries between the tumor and the surrounding tissue are often unclear, the marking process is essential for visually distinguishing the boundaries of the tumor [46]. This is also used to plan the cutting trajectory, so it is preferable to define the marking points in advance before the cutting process. Therefore, this process should be carried out even in the robotic ESD automation process to generate the primitive cutting motion trajectory and the safety of the endoscopic surgery.
@@
In robotic ESD automation for tumor removal, safety may be compromised due to an abnormal reference trajectory for the robot or unexpected human-induced errors during teleoperation; thus, it is crucial to consider safety guarantees. For example, the control design for the surgery automation should prevent the endoscopic robot from damaging tumors. As damaging a tumor can alter the cells or tissues inside it, it is crucial to cut the tumor without causing any damage for accurate examination and pathological analysis.
@@
$\mathcal{S}_{n}=\{\bm{x}\in\mathcal{X}\text{ }|\text{ }||\bm{x}-P_{n}^{\top}||\geq r_{n}\},$ (20)
@@ III-B Robust Model-Free Safety Critical Control
where $D=||\bm{x}-P_{n}^{\top}||$ is the distance between the current end-effector position and the center point of the tumor $P_{n}$. The gradient of the safe set, $\nabla h_{n}(\bm{x})=\frac{(\bm{x}-P_{n}^{\top})^{\top}}{||\bm{x}-P_{n}^{\top}||}$, is equal to the unit vector $\mathsf{n}_{n}^{\top}$, which points from the tumor center to the endoscopic robot. Lastly, the safe velocity controller used in [40] is designed by formulating the following quadratic program that minimally deviates from the desired velocity $\dot{\bm{x}}_{d}$:
@@
where $\dot{\bm{e}}=J^{+}(\bm{q})(\dot{\bm{x}}-\dot{\bm{x}}_{s})$, and $J^{+}(\bm{q})$ is the pseudo-inverse of the manipulator Jacobian matrix, $J(\bm{q})$, and $k_{d}\in\mathbb{R}_{\geq 0}$ is the controller gain. The control law, (24) does not include the model dynamics terms, $M(\bm{q}),C(\bm{q},\dot{\bm{q}}),g(\bm{q})$, but the controller satisfies the ISS condition from (17). From Corollary 1, we ensure that $\mathcal{S}^{*}$ is forward invariant with the velocity controller, (24), tracking $\dot{\bm{x}}_{s}$, which ensures the ISSf condition with respect to $\mathcal{S}^{*}$. It is worth noting that, to the best of our knowledge, this is the first time that formal safety guarantees in robotic ESD automation using a well-known model-free control law such as (24), are explored.
@@
IV Experimental Design
@@ IV-A Mechanical Design and Numerical Simulation of an Endoscopic Robot Manipulator
A three degree-of-freedom (DOF) endoscopic robot manipulator was designed to perform cutting automation for ESD in a three-dimensional space (Fig. 4). This manipulator consists of one translational DOF and two rotational DOF. From the previous studies on the mechanical design for surgical robots with endoscopes [12, 16, 17], the kinematic parameters of the endoscopic robot manipulator used in this experiment, are decided as follows: $l_{1}$ = 3 mm, $l_{2}$ = 10 mm, $l_{end}$ = 17 mm, $OD$ = 3.7 mm.
@@ IV-B Experiment Scenarios
Fig. 5 shows an overview of the experimental scenarios performed in this study. A maximum of two tumors exist within the task environment of the endoscopic robot, each was defined as either a tumor that needs to be removed or one that should be preserved. In this study, three scenarios were defined for this experiment as follows: (Scenario-1) A single tumor with two unsafe marking points inside the safe set boundary (Fig. 5(a)); (Scenario-2) Two tumors in close proximity, with one unsafe marking point inside the safe set boundary of the tumor that should not be removed (Fig. 5(b)); (Scenario-3) Two tumors in close proximity, with two additional unsafe marking points inside the safe set boundary of the tumor that needs to be removed (Fig. 5(c)). The green-filled circles represent ideal safe marking points, while the red-filled circles represent unsafe marking points that may pose clinical issues. Additionally, the black dashed circles represent the safe set boundary of the tumor that needs to be removed, while the blue dashed circles represent the safe set boundary of the tumor that should not be removed. To generate the reference trajectory that robot should track, we linearly interpolate the marking points including safe and unsafe points.
@@
Figure 6: The simulation results of the proposed method. We conduct three scenarios with unsafe marking points. The rows show the results of each scenario and the three columns correspond to the position of the end-effector, safety guarantees with $h(\bm{x})\geq 0$, the performance of velocity tracking controller with $\alpha=0.4$, respectively.
@@ V Results
In the first scenario, as shown in the first row of Fig. 6, the end-effector position is trying to track the reference trajectory, $P_{ref}$ defined by the marking points. Even if the reference trajectory would cause the robot to cross into the tumor due to the unsafe marking points (red circle), the robot still achieves safe cutting motions without crossing the safety boundary (dashed green line) with the help of the CBF-based safe controller (23). It is clearly observed that the safety guarantees are enforced at all times as, shown in the safe constraint function, $h(\bm{x})$ of Fig. 6.
Furthermore, several simulations are carried out by varying the parameter, $\alpha$, in order to evaluate the level of conservatism. As shown in the middle graph of the first row of Fig. 6, it is observed that the conservatism decreases as the alpha value increases. Note, however, that the safety guarantees might be violated if we choose an excessively large alpha [40]. Next, it is also shown that the velocity controller accurately tracks the safe velocity provided by the CBF-based safe controller (23) in the rightmost graph in the first row. It is worth noting that the CBF-based safe controller (23) provides the control input that is closest to the desired velocity reference, while at the same time ensuring safety constraints.
In the second and third graphs, the safety guarantees with the proposed framework are still achieved like the first scenario, although there exist unsafe marking points in the tumors as shown in the second and third rows of Fig. 6. Depending on $\alpha$, similar conservative behavior to that observed in the first scenario can be seen in these two scenarios, as shown in the middle graphs of the second and third rows. Likewise, it is observed that the controller tracks the safe velocities in these two scenarios with results similar to the first scenario.

### v2

@@ Abstract
There is growing interest in automating surgical tasks using robotic systems, such as endoscopy for treating gastrointestinal (GI) cancer. However, previous studies have primarily focused on detecting and analyzing objects or robots, with limited attention to ensuring safety, which is critical for clinical applications, where accidents can be caused by unsafe robot motions. In this study, we propose a new control framework that can formally ensure the safety of automating the cutting task in endoscopic submucosal dissection (ESD), a representative endoscopic surgical method for the treatment of early GI cancer, by using an endoscopic robot. The proposed framework utilizes Control Barrier Functions (CBFs) to accurately identify the boundaries of individual tumors, even in close proximity within the GI tract, ensuring precise treatment and removal while preserving the surrounding normal tissue. Additionally, by adopting a model-free control scheme, safety assurance is made possible even in endoscopic robotic systems where dynamic modeling is challenging. We demonstrate the proposed framework in a simulation-based experimental environment, where the tumors to be removed are close to each other, and show that the safety constraints are enforced. We show that the model-free CBF-based controlled robot eliminates one tumor completely without damaging it, while not invading another nearby tumor.
@@ I Introduction
One of the most widespread methods to ensure safety in the robot movement is virtual fixtures. Virtual fixture is a method that defines virtual boundaries or constraints to guide the movements of a robot along a specific trajectory or restrict its operation within a designated space [30]. Especially in the field of surgical robotics, it has been widely applied to improve safety by preventing the robot from approaching sensitive organs or lesions [31]. By using both guidance virtual fixtures and forbidden region virtual fixtures together, the surgical robot can track an accurate trajectory while avoiding areas that should not be accessed. However, the mathematical expressions that define virtual fixture constraints typically exhibit discontinuities. This is an issue, because such discontinuous constraints can cause damage to the lesion tissue due to unintended movements, significantly impacting the quality and reliability of robotic surgery. Furthermore, virtual fixture-based approaches lack formal safety guarantees, which might lead to serious safety violations.
@@
Based on Molnar’s work [40], we propose a new framework that can formally ensure the safety of certain processes involved in endoscopic submucosal dissection (ESD), a representative endoscopic surgical method for the treatment of early GI cancer [1, 2, 42], by using an endoscopic robot (Fig. 1). The purpose of ESD is not only to completely remove the tumor from the GI tract but also to ensure that the entire tumor is resected as a single piece, allowing for accurate pathological examination. In addition to the tumor boundaries, it is critical that the cutting depth is shallow enough to only pierce the submucosa, as too deep a cut could result in bleeding or puncturing of the intestine. In summary, as shown in Fig. 2, the framework proposed in this study enables the automated motion of the robot to clearly respect the boundaries of each tumor, even when one or more tumors are located close to each other within the GI tract, while enforcing depth constraints. This allows for the individual removal of each tumor while preserving the normal mucosal areas between tumors.
The main contribution of this study can be summarized as follows:
  • This paper proposes an integrated framework that ensures the safe automation of the cutting task in ESD by fusing clinical and engineering perspectives. The framework plans safe motions for tumor removal while enforcing safety boundaries as constraints to prevent unintended damage to surrounding tissue.
  • This study extends the existing model-free CBF to medical robotic applications in where we consider clinical safety in the depth of the cut, contributing to prevent complications such as gastrointestinal perforation.
@@ II-C Input-to-State Stability
where the bounded disturbance, $\bm{d}\in\mathbb{R}^{m}$, deteriorates the performance of CLF in terms of stability. To resolve the impact of disturbances, the concept of Input-to-State Stability (ISS) [43] can be introduced in (3) as follows:
@@
where $\iota$ is an extended class $\mathcal{K}^{e}_{\infty}$ function; thus, we can achieve a robust stable controller satisfying (7). It is worth noting that the concept allows the controller to converge to a neighborhood of the origin, is case of a nonzero disturbance. $\iota$ can be determined based on how conservatively the system should operate with respect to disturbances [43].
@@ II-E Model-Free Robotic Safety-Critical Control
There exists the exponentially stable velocity controller, $\bm{u}=k_{q}(\bm{q},\dot{\bm{q}})$ where $k_{q}:\mathrm{Q}\times\mathbb{R}^{n}\rightarrow\mathcal{U}$, satisfying $||\dot{\bm{e}}(t)||\leq M||\dot{\bm{e}}_{0}||\mathrm{e}^{-\lambda t}$ for the positive scalar, $M,\lambda$, where $\dot{\bm{e}}=\dot{\bm{q}}-\dot{\bm{q}}_{d}$ is the tracking error between current velocity and desired one. In other words, a continuously differentiable Lyapunov function, $V(\bm{q},\dot{\bm{e}})$ exists such that:
$k_{1}||\dot{\bm{e}}||\leq V(\bm{q},\dot{\bm{e}})\leq k_{2}||\dot{\bm{e}}||,\quad k_{1},k_{2}\in\mathbb{R}_{>0}$ (12)
@@
and a stable velocity tracking controller satisfying (13). If the velocity controller tracks a safe velocity, $\dot{\bm{q}}_{s}$, fast enough [40], and if $\lambda>\alpha$ such that the initial conditions, $(\bm{q}_{0},\dot{\bm{e}}_{0})\in\mathcal{S}_{V}$ can be defined as:
@@
Figure 2: Illustration of the proposed framework for safe robotic ESD cutting task automation. Once the tumor is detected by the robot, the safe set boundary is defined to ensure safety. Then, a series of marking points are automatically generated around the tumor. Finally, cutting is performed along the trajectory defined by these marking points using a CBF-based safety-critical velocity controller.
III Application to Safety-Critical Robotic Cutting Task Automation in ESD
Fig. 2 illustrates the framework we propose for ensuring formal safety in ESD cutting task automation. ESD can be subdivided into four main processes: injection, marking, cutting, and dissection [46]. Considering the difficulty of each process involved in ESD and the potential clinical safety issues, this study focuses on developing a method to ensure the safety of marking and cutting automation.
@@ III-A Automatic Definition of Marking Points
The marking process is to mark specific points around each tumor. Since the boundaries between the tumor and the surrounding tissue are often unclear, the marking process is essential for visually distinguishing the boundaries of the tumor [46]. This is also used to plan the cutting trajectory, so it is preferable to define the marking points in advance before the cutting process. Therefore, this process should be carried out even in the robotic ESD cutting task automation process to generate the primitive cutting motion trajectory and the safety of the endoscopic surgery.
@@
In robotic ESD cutting task automation for tumor removal, safety may be compromised due to an abnormal reference trajectory for the robot or unexpected human-induced errors during teleoperation; thus, it is crucial to consider safety guarantees. For example, the control design for the surgery automation should prevent the endoscopic robot from damaging tumors. As damaging a tumor can alter the cells or tissues inside it, it is crucial to cut the tumor without causing any damage for accurate examination and pathological analysis. Furthermore, as too deep a cut can result in bleeding or damage to the intestine, it is critical for the control framework to also constrain the allowed depth.
@@
$\mathcal{S}_{n}=\{\bm{x}\in\mathcal{X}\text{ }|\text{ }||\bm{x}-P_{n}||\geq r_{n}\},$ (20)
@@ III-B Robust Model-Free Safety Critical Control
where $D=||\bm{x}-P_{n}||$ is the distance between the current end-effector position and the center point of the tumor $P_{n}$. The gradient of the safe set, $\nabla h_{n}(\bm{x})=\frac{(\bm{x}-P_{n})^{\top}}{||\bm{x}-P_{n}||}$, is equal to the unit vector $\mathsf{n}_{n}^{\top}$, which points from the tumor center to the endoscopic robot. Lastly, the safe velocity controller used in [40] is designed by formulating the following quadratic program that minimally deviates from the desired velocity $\dot{\bm{x}}_{d}$:
@@
where $\dot{\bm{e}}=J^{+}(\bm{q})(\dot{\bm{x}}-\dot{\bm{x}}_{s})$, and $J^{+}(\bm{q})$ is the pseudo-inverse of the manipulator Jacobian matrix, $J(\bm{q})$, and $k_{d}\in\mathbb{R}_{\geq 0}$ is the controller gain. The control law, (24) does not include the model dynamics terms, $M(\bm{q}),C(\bm{q},\dot{\bm{q}}),g(\bm{q})$, but the controller satisfies the ISS condition from (17). From Corollary 1, we ensure that $\mathcal{S}^{*}$ is forward invariant with the velocity controller, (24), tracking $\dot{\bm{x}}_{s}$, which ensures the ISSf condition with respect to $\mathcal{S}^{*}$. It is worth noting that, to the best of our knowledge, this is the first time that formal safety guarantees in robotic ESD cutting task automation using a well-known model-free control law such as (24), are explored.
@@
IV Simulation-Based Experimental Design
@@ IV-A Mechanical Design and Numerical Simulation of an Endoscopic Robot Manipulator
A three-degree-of-freedom (DOF) endoscopic robot manipulator was designed to perform cutting automation for ESD in a three-dimensional space (Fig. 4). This manipulator consists of one translational DOF and two rotational DOFs. To define the key parameters, such as the DOFs, outer diameter (OD), and total length, we referred to the geometric information of robots designed in previous studies on real endoscopic robot systems [12, 16, 17]. Based on these studies, the kinematic parameters of the endoscopic robot manipulator used in this experiment are as follows: $l_{1}$ = 3 mm, $l_{2}$ = 10 mm, $l_{end}$ = 17 mm, $OD$ = 3.7 mm.
@@ IV-B Experiment Scenarios
Fig. 5 shows an overview of the experimental scenarios performed in this study. A maximum of two tumors exist within the task environment of the endoscopic robot, each was defined as either a tumor that needs to be removed or one that should be preserved. In this study, four scenarios were defined for this experiment as follows: (Scenario-1) A single tumor with two unsafe marking points inside the safe set boundary (Fig. 5(a)); (Scenario-2) Two tumors in close proximity, with one unsafe marking point inside the safe set boundary of the tumor that should not be removed (Fig. 5(b)); (Scenario-3) Two tumors in close proximity, with two additional unsafe marking points inside the safe set boundary of the tumor that needs to be removed (Fig. 5(c)); (Scenario-4) An additional depth constraint for safety is considered above the third scenario (Fig. 7). The blue-filled circles represent ideal safe marking points, while the red-filled circles represent unsafe marking points that may pose clinical issues. Additionally, the green dashed circles represent the safe set boundary of each tumor. To generate the reference trajectory that robot should track, we linearly interpolate the marking points including safe and unsafe points. The final reference trajectory is colored as red dashed lines.
@@
Figure 6: The simulation results of the proposed method. We conduct three scenarios with unsafe marking points. The rows show the results of each scenario and the three columns correspond to the position of the end-effector, safety guarantees with $h(\bm{x})\geq 0$, the performance of velocity tracking controller with $\alpha=0.4$, respectively. Note that $V_{d}$, $V_{safe}$, and $V_{curr}$ are the desired, safe, and current velocities, respectively.
@@ V Results
Figure 7: The simulation results of the scenario where we take into account depth violations. We use the same reference trajectory as what used in the third scenario. The first graph shows the whole position trajectory of the end-effector depending on $\alpha$. The second shows safety guarantees with $h(\bm{x})\geq 0$, and lastly the safe velocity tracking with $\alpha=1.5$ is shown in the third graph. Note that the safe controller is activated after the robot reaches the safe area, which happens about $1$ second.
@@
In the first scenario, as shown in the first row of Fig. 6, the end-effector position is trying to track the reference trajectory, $P_{ref}$ defined by the marking points. Even if the reference trajectory would cause the robot to cross into the tumor due to the unsafe marking points (red circle), the robot still achieves safe cutting motions without crossing the safety boundary (dashed green line) with the help of the CBF-based safe controller (23). It is clearly observed that the safety guarantees are enforced at all times as shown in the safe constraint function, $h(\bm{x})$ of Fig. 6.
Furthermore, several simulations are carried out by varying the parameter, $\alpha$, in order to evaluate the level of conservatism. As shown in the middle graph of the first row of Fig. 6, it is observed that the conservatism decreases as the alpha value increases. Note, however, that, as will be discussed in the following section, the safety guarantees might be violated if we choose an excessively large alpha [40]. Next, it is also shown that the velocity controller accurately tracks the safe velocity provided by the CBF-based safe controller (23) in the rightmost graph in the first row. It is worth noting that the CBF-based safe controller (23) provides the control input that is closest to the desired velocity reference, while at the same time ensuring safety constraints.
In the graphs of the second and third rows, the safety guarantees with the proposed framework are still achieved like the first scenario, although there exist unsafe marking points in the tumors as shown in the second and third rows of Fig. 6. Depending on $\alpha$, similar conservative behavior to that observed in the first scenario can be seen in these two scenarios, as shown in the middle graphs of the second and third rows. Likewise, it is observed that the controller tracks the safe velocities in these two scenarios with results similar to the first scenario.
To demonstrate that our proposed framework can also be used ensure that the depth of the cut is kept shallow enough to avoid damage to the intestine or submucosal tissue, we conduct an additional scenario where we include depth constraint while cutting a tumor and also compare the results with different parameters, $\alpha$. As shown in Fig. 7, we defined a larger outer constraint sphere covering the tumor, colored pale green, in order to consider cutting depth. Subsequently, we designed the safe set between the inner sphere (a tumor to be removed) and the outer one while using (22) and the control barrier function as defined by $h_{n}(\bm{x})=-D+r_{n}\geq 0$. As discussed in Molnar et al. [40], the closest control barrier function is considered to provide the safe velocity reference in equation (23). For example, when the robot is about to enter the tumor, the direction of the safe velocity, $\frac{(\bm{x}-P_{n})^{\top}}{||\bm{x}-P_{n}||}$ should be outward. Conversely, when the robot approaches the boundary of the outer sphere, the direction of the safe velocity should be inward. As a result, the safe cutting trajectory is also ensured with depth constraint as shown in the first and second graphs of Fig. 7, while showing that $h(\bm{x})$ is always over 00. Likewise, it is found that the larger $\alpha$ reduces the conservatism of the controller as well. Compared to the previous scenarios, we choose a higher $\alpha$ value since the current case has the smaller safe region, and the less conservative performance of the safe controller is necessary to ensure the safety guarantees with the region. Note that $\alpha$ values used in the simulations were arbitrarily selected to ensure the safety guarantees. Tuning the values requires additional analysis, which is beyond the scope of this paper; we will briefly discuss this aspect in the following section.